If 2×10^9 electrons are transferred from a glass rod to a piece of silk by rubbing,what will be the net charge on both objects?
1
Expert's answer
2021-08-23T12:40:37-0400

Electron Transferred from glass to the silk "= 2 \times 10^{9}"

Charge on one electron "=-1.602\times 10^{-19} \; C"

Hence total charge transferred from Glass to silk

"=(-1.602\times 10^{-19} \; C) \times (2.0\times 10^{9}) \\

=-3.204 \times 10^{-10} \;C"

Because "-3.204 \times 10^{-10} \; C" charge is transferred from glass to silk, hence the silk will have a net negative charge and glass will have a net positive charge values of which are given below:

Net charge on silk "=-3.204 \times 10^{-10} \; C"

Net charge on Glass "=3.204 \times 10^{-10} \; C"